The difference of the squares of two numbers is 10. The difference of the two numbers is 2. What is their sum?
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

The difference of the squares of two numbers is 10.
The difference of the two numbers is 2. What is their sum?
---------------
x^2-y^2 = 10
x-y = 2
---
Divide the 1st equation by the 2nd equation to get:
x+y = 5
----
Solve the following system:
x-y = 2
x+y = 5
----------
Add and solve for "x":
2x = 7
x = 7/2
---------
Substitute for "x" and solve for "y":
x + y = 5
(7/2) + y = 5
y = (10/2)-(7/2)
y = 3/2
===========
Cheers,
Stan H.
===========

Answer by richwmiller(17219)

x^2-y^2=10
x-y=2
(x+y)*(x-y)=10
(x+y)*2=10
x+y=5
x-y=2
2x=7
x=3.5
3.5-y=2
y=1.5